# Global-Local Progressive Integration Network for Blind Image Quality Assessment

### v1

@@ II-A Architecture of BIQA Models
Figure 1: The framework of the proposed GlintIQA. The input image is processed by both VGFE and CLFE for global and local feature extraction, followed by feature alignment and progressive integration through interactive stacking of CWSA and SEM. The resulting multi-grained representation is fed to an MLP for image quality prediction. For clarity, dashed lines represent omitted network layers preceding CLFE and VGFE.
II-B Data Augmentation of BIQA Models
Regarding the data aspect, researchers addressed overfitting issues in size-restricted IQA databases through augmentation techniques and training strategies, including patch-based augmentation methods and pre-training with expanded datasets. The patch-based approach randomly selects image patches of varying sizes, such as 32$\times$32 employed in [7] or 256$\times$256 employed in [9], from the full images. The corresponding mean opinion scores of these patches then serve as the regression targets for model training. To mitigate potential label bias from small patches, Kim et al.[35] utilized patch-level quality scores from FR-IQA methods as intermediate regression targets. Furthermore, several studies [9, 10, 27, 36, 28] created large-scale datasets for pre-training BIQA models. For instance, Ma et al.[9] synthesized images with different distortions using artificial degradation for training distortion classification networks. Liu et al.[10] leveraged injected distortion levels of synthetic images as quality rankings. Ma et al.[27] and Gu et al.[36] calculated quality scores for a large number of synthetically distorted images using either a single FR-IQA method [37] or by fusing the results of multiple FR-IQA methods in [28].
Previous work has explored data augmentation strategies for synthetic distortions, using injected distortion levels as quality labels or FR-IQA algorithms for quality annotation. However, these methods have inherent shortcomings. Distortion level labels can only partially reflect image quality that cannot capture the continuous degradation of image quality. Although FR-IQA algorithms can generate fine-grained image quality scores, they may exhibit biases towards certain distortion types. For instance, while Peak Signal-to-Noise Ratio (PSNR) adequately estimates images corrupted by additive noise and quantization artifacts, it falls short in assessing numerous other distortion types [25]. In addition, these FR-IQA algorithms often rely on low-level feature discrepancies to quantify distortion severity, overlooking the significant role of high-level semantic content in visual quality perception.
@@ III The Proposed GlintIQA
As illustrated in Fig. 1 The proposed GlintIQA first leverages a hybrid feature extraction scheme combining ViT-based global feature extractor and CNNs-based local feature extractor, which respectively leverages ViT for capturing global coarse-grained features and CNNs for extracting fine-grained details at multiple scales. Then, interactive stacks of channel-wise attention and local spatial enhancement modules progressively integrate these features to build multi-grained representations. The integrated multi-grained features are finally passed through an MLP to generate the perceptual image quality score.
@@ III-A Global and Local Feature Extraction
III-A1 ViT-based Global feature extractor (VGFE)
@@
$\textbf{f}_{g}^{e}=\psi_{\tiny f}\left(\varphi_{conv(3,d,k)}\left(\textbf{x}\right)\right)^{\text{T}},$ (1)
where the notation $\varphi_{conv(3,d,k)}$ represents a 2D convolutional layer with convolutional kernel size and stride of $k$ that embeds the input image with 3 channels into a feature space of $d$ dimensions. The function ${{\Large\psi}}_{\tiny f}$ represents the flattening operation that collapses the output of the convolutional layer along the spatial dimensions, yielding $\textbf{f}_{g}^{e}\in\mathbb{R}^{d\times\frac{HW}{k^{2}}}$. Subsequently, $\textbf{f}_{g}^{e}$ is transposed to obtain features within $\mathbb{R}^{\frac{HW}{k^{2}}\times d}$, where each patch of size $k\times k$ is represented by a vector of length $d$, referred to as a token. For simplicity, we have omitted the description of the class token and positional embedding used in the original ViT architecture.
$\textbf{f}_{g}^{i}=\mathcal{E}_{g}^{i}\left(\textbf{x};\omega_{g}^{i}\right),$ (2)
Although the patch embedding process, as shown in Eq. 1, enables subsequent models to capture global context, it may result in the loss of fine-grained details within patches. As shown in Fig. 2(a), patch embedding discretizes the input image into a sequence of visual tokens, significantly reducing spatial dimensions in a single step at the model input. This discretization inevitably leads to a loss of fine-grained details and local spatial information, as the individual tokens correspond to low-resolution patches rather than individual pixels. To address this limitation, we exploit a local feature extractor consisting of a series of convolutional layers, as depicted in Fig. 2(b). This extractor uses small convolutional kernels and strides to progressively reduce the dimensions of the image, which can preserve certain local image details in the extracted features. Moreover, the integrated CNNs-based structure can mitigate the inherent lack of inductive bias in ViT. By leveraging the strengths of both CNNs and ViTs, the hybrid model effectively captures both global and local features, ensuring a more accurate and robust representation of image quality.
@@
III-A2 CNNs-based Local feature extractor (CLFE)
Local features, such as edges and textures, reside within small image regions and provide valuable insights into the nature and severity of local distortions. To meticulously extract these features, the CLFE utilizes three residual blocks from ResNet50 [38], denoted as $\mathcal{E}_{l}^{j}$ with parameters $\omega_{l}^{j}$, where $j$ represents the index of the blocks. Each block yields features at distinct scales, downsampled by factors of $2^{j+1}$ relative to the original image.
@@ III-A2 CNNs-based Local feature extractor (CLFE)
$\textbf{f}_{l}^{j}={{\Large\psi}}_{\tiny f}\left(\varphi_{conv(c_{j},d,k/{2^{j+1}})}\left({\mathcal{E}}_{l}^{j}\left(\textbf{x};\omega_{l}^{j}\right)\right)\right)^{\text{T}},$ (3)
@@
$\textbf{f}_{l}=\varphi_{p(3d,d)}\left(\mathcal{M}_{\text{CWSA}}\left(\textbf{f}_{l}^{j-1}\oplus\textbf{f}_{l}^{j}\oplus\textbf{f}_{l}^{j+1}\right)\right),$ (4)
where $\textbf{f}_{l}\in\mathbb{R}^{\frac{HW}{K^{2}}\times d}$ and $\oplus$ denotes a concatenation operation, and $\varphi_{p(3d,d)}$ denotes a linear projection layer with input and output dimensions of $3d$ and $d$, respectively. The CWSA module, denoted as $\mathcal{M}_{\text{CWSA}}$, is described in detail in the following subsection.
(a) The structure of CWSA
(b) The structure of SEM
Figure 3: Structures of (a) channel-wise Self-attention (CWSA) and (b) Spatial Enhancement Module (SEM).
III-B Local and Global Feature Integration
@@ III-B1 Progressive Feature Integration
Directly concatenating and fusing features at multiple granularities may pose challenges for the model to handle diverse information simultaneously. Therefore, we adopt a progressive approach to integrate features layer by layer while optimizing the channel and spatial representations of the fused features. Specifically, we start by concatenating the fine-grained feature $\textbf{f}_{l}$ with the coarse-grained global feature at the initial scale. Then, a stacked combination of CWSA and a SEM is used to progressively fuse features in a hierarchical manner. The feature integration process can be concisely represented as:
$\textbf{z}_{i}$ $=\left(\textbf{f}_{i}\oplus\textbf{f}_{g}^{i}\right)^{\text{T}};$ (5)
@@
$\hat{\textbf{z}}_{i}$ $=\mathcal{M}_{\text{CWSA}}\left(\textbf{z}_{i}\right);$
$\textbf{f}_{i+1}$ $=\mathcal{M}_{\text{SEM}}\left(\hat{\textbf{z}}_{i}\right).$
@@
For $i=1$, $\mathbf{f}_{i}$ represents the output feature of the CLFE, i.e., $\mathbf{f}_{l}$. For other values of $i$, $\mathbf{f}_{i}$ is the output feature of the SEM, denoted as $\mathcal{M}_{\text{SEM}}$.
III-B2 Channel-wise Self-attention (CWSA)
This module is designed to learn the importance of different channels in features (extracted by CNNs and ViTs). By accurately weighting these channels, the model can emphasize the most informative features [39], leading to improved overall performance. As illustrated in Fig. 3(a), CWSA concatenates two inputs to obtain the feature $\textbf{z}_{i}$ as shown in Eq. 5. Subsequently, the feature is transposed and subjected to a self-attention computation. Specifically, $\textbf{z}_{i}$ is projected into three representations for query, key, and value through three distinct linear layers, which is expressed as:
$\textbf{Q}=\varphi^{Q}_{p(d,d)}(\textbf{z}_{i});~{}\textbf{K}=\varphi^{K}_{p(d,d)}(\textbf{z}_{i});~{}\textbf{V}=\varphi^{V}_{p(d,d)}(\textbf{z}_{i}).\$ (6)
The softmax scores quantify the channel-wise significance, which get applied to the scaled dot product of queries (Q) and keys (K), normalized by $\sqrt{d}$ ($d$ being the dimension). The CWSA output is the weighted summation of values (V) based on these attention scores, combined with the previous input feature $\textbf{z}_{i}$, which is expressed as:
$\hat{\textbf{z}}_{i}=\varphi_{\text{Softmax}}(\text{Q}\cdot\text{K}^{\text{T}}/\sqrt{d})\cdot\text{V}+\textbf{z}_{i}.$ (7)
III-B3 Spatial Enhancement Module (SEM)
Following the CWSA module, we introduce SEM to enhance the interaction of features extracted by CNNs and ViT. The SEM reshapes and applies convolutional operations to the tokens for the input tensor $\hat{\textbf{z}}_{i}$, as illustrated in Fig. 3(b). The computational process is expressed as:
$\textbf{f}_{i+1}=\psi_{f}\left(\varphi_{conv(2d,d,3)}\left(\psi_{r}\left(\hat{\textbf{z}}_{i}\right)\right)\right),$ (8)
@@ III-B3 Spatial Enhancement Module (SEM)
where the notation ${\Large\psi}_{\tiny r}$ indicates the operation of reshaping the token vectors into 2D feature maps. The convolutional layer $\varphi_{conv(2d,d,3)}$ employs a $3\times 3$ filter with a stride of 1. Finally, the output of the convolutional layer is reshaped back to a flattened tensor, aligning it with the original input dimensions.
III-C Image Quality Prediction.
@@
The quality score prediction is performed via an MLP architecture, which consists of two linear layers interspersed with GELU activation function and dropout layers for regularization. Specifically, the feature map output from the SEM module is denoted as f. To generalize the global distribution of these features and obtain a concise quality representation, feature $\text{f}_{5}$ is spatially averaged to form a $d$-dimensional vector $\hat{\textbf{f}}\in\mathbb{R}^{1\times d}$. This is fed into the MLP to generate a prediction score $s$, which can be represented as:
$\text{s}=\varphi_{p(d,1)}\left(\varphi_{p(d,d)}(\hat{\textbf{f}})\right).$ (9)
@@
For simplicity, the specific details of the activation function and dropout layer are omitted from the above equation. Ultimately, $l_{1}$ loss function is utilized to enable end-to-end model parameter updates.
Figure 4: Analyzing workflow of correlation between image content similarity and visual perception quality.
IV The proposed content similarity-guided labeling method
Deep neural networks typically possess a large number of learnable parameters, necessitating access to large-scale and diverse training datasets to facilitate effective learning and achieve better generalization capability. However, synthetic distortion datasets often lack content diversity, necessitating effective data augmentation. Patch-based augmentation method, while expanding the training set, introduces labeling biases that hinder performance improvements. Furthermore, FR-IQA-labeled data may suffer from prediction biases for specific distortion types and fail to capture high-level semantic impacts on quality. To address these limitations, this section introduces a novel annotation strategy based on quality label transfer using content similarity. We first evaluate the effectiveness of this approach in automatically annotating image quality, followed by a detailed explanation of the construction process for a large-scale synthetic distortion dataset.
IV-A MOS Correlation Analysis for Semantically Similar Images
The proposed labeling method draws inspiration from empirical findings that images with similar content tend to have comparable visual quality when degraded by the same type and level of distortion [40]. To further substantiate the viability of this method, we conduct an analysis of an example using the KADID-10k [24] as shown in Fig. 4. Specifically, given a pritine image $X_{i}$ and its corresponding distorted images $\textbf{X}_{i,d}=\left\{x^{1,1}_{i,d},…,x^{k,l}_{i,d}\right\}$, where $k$ and $l$ respectively represent the $k$-th type and $l$-th level of distortion, we utilize semantic feature vector to represent the image content. Then, a pristine image $X_{j}$ with similar content to $X_{i}$ was retrieved using cosine similarity between semantic vectors. The MOS values of distorted images $\textbf{X}{j,d}$ from KADID-10k were then used as predicted quality scores for $\textbf{X}{i,d}$. To benchmark this evaluation, we analyzed the results for all pristine images in KADID-10k and compared it against several FR-IQA algorithms, i.e., SSIM [41], FSIMc [42], and VSI [37]. As illustrated in Fig. 5, our method outperforms these FR-IQA algorithms in the majority of cases, as evidenced by the SROCC and RMSE metrics. In a comprehensive evaluation of 81 distorted image sets, our method achieved the best SROCC in 88.9% (72 of the sets) and the best RMSE in 86.4% (70 of the sets), which demonstrate superior correlation and accuracy values, respectively. Hence, images with similar content exhibit similar visual quality labels under identical or similar distortion conditions. Based on this idea, the proposed approach involves leveraging human-annotated images to identify content-similar images, applying the same distortions, and subsequently transferring the annotations. Next, we describe the specific steps for constructing the proposed content similarity-guided dataset.
(a) Correlation
(b) Accuracy
Figure 5: The results of our labeling method. Our labeling method achieved superior performance compared to these FR-IQA methods in the majority of cases for correlation and accuracy metrics, when evaluating distorted versions of 81 pristine images in KADID-10k.
Figure 6: The scheme of the proposed content similarity-guided labeling method.
IV-B Dataset Construction
IV-B1 Preliminary.
To create the dataset, we selected 50,000 high-quality images from KADIS-700k [24] and applied 25 types of distortions with five intensity levels to each image. The distortion types were selected to be consistent with those present in KADID-10k [24]. Finally, this process produced 6,250,000 degraded images.
The quality of the generated images was annotated using MOS values obtained from an IQA database, i.e., KADID-10k, where the IQA database with $P$ pristine images is denoted as $\left\{(\textbf{X}_{i,d},\textbf{M}_{i,d})\right\}_{i=1}^{P}$. The set $\textbf{X}_{i,d}=\left\{x^{1,1}_{i,d},…,x^{k,l}_{i,d}\right\}$ represents the degraded images of the $i$-th pristine image $X_{i}$, and $\textbf{M}_{i,d}=\left\{m^{1,1}_{i,d},…,m^{k,l}_{i,d}\right\}$ contains their corresponding MOS values, which were normalized using min-max scaling and mapped to a scale ranging from 0 to 9. Each degraded image $x^{k,l}_{i,d}$ corresponds to a pristine image $X_{i}$ degraded with the $k$-th type at the $l$-th level of distortion, with its MOS value given by $m^{k,l}_{i,d}$. The high-quality images are represented as a set denoted by $\left\{I_{j}\right\}_{j=1}^{H}$, where $I_{j}$ and $H$ refer to the $j$-th high-quality image and the total number of images, respectively.
IV-B2 Dataset construction.
The process of constructing the dataset is shown in Fig. 6. To start the process, the first objective is to retrieve a pristine image $X_{i}$ that exhibits similar content to a given image $I_{j}$. To achieve this, we extract the semantic feature vectors of $X_{i}$ and $I_{j}$ using ResNet101 [38] without the final linear layer, where these semantic vectors denote as $\textbf{v}_{i}$ and $\textbf{v}_{j}\in\mathbb{R}^{1\times 2048}$, respectively. Following that, the similarity between their image contents is assessed by calculating the cosine similarity between their semantic features. The most similar pristine image is identified by:
$\hat{i}=\mathop{\mathrm{argmax}}\limits_{i\in\left\{1,...,P\right\}}\cos(\textbf{v}_{i},\textbf{v}_{j})=\frac{\textbf{v}_{i}\cdot\textbf{v}_{j}^{T}}{\|\textbf{v}_{i}\|\cdot\|\textbf{v}_{j}\|}.$ (10)
In the second step, based on the retrieved pristine image $X_{\hat{i}}$, we obtain corresponding distorted images $\textbf{X}_{\hat{i},d}$ and MOS values $\textbf{M}_{\hat{i},d}$ from IQA database. The third step involves degrading $I_{j}$ using the same type and level of distortion as applied to the pristine image $X_{\hat{i}}$. This degradation process is represented by an operation $W^{k,l}_{\hat{i}}(\cdot)$ that generates a degraded image denoted as $I^{k,l}_{j,d}$ [^2] The synthetic distortion procedure available at: here.. In the fourth step, the MOS value $m^{k,l}_{\hat{i},d}$ corresponding to $x^{k,l}_{\hat{i},d}$ is used to annotate the degraded image $I^{k,l}_{j,d}$. The generation of degraded images and the labeling process can be described as
$\left\{(I^{k,l}_{j,d},s^{k,l}_{j,d})~{}|~{}I^{k,l}_{j,d}=W^{k,l}_{\hat{i}}(I_{j}),s^{k,l}_{j,d}=m^{k,l}_{\hat{i},d},1\leq j\leq H\right\}.$ (11)
Finally, the generated dataset can be represented as $\left\{(\textbf{I}_{j,d},\textbf{S}_{j,d})\right\}_{j=1}^{H}$, where $\textbf{I}_{j,d}=\left\{I^{1,1}_{j,d},…,I^{k,l}_{j,d}\right\}$ and $\textbf{S}_{j,d}=\left\{s^{1,1}_{j,d},…,s^{k,l}_{j,d}\right\}$. The proposed content-guided label transfer method generates quality scores closely aligning with the true distribution of subjective opinion scores, as labels derive from genuine human annotations. Moreover, the proposed approach considers image content influence on quality perception in a content-driven manner. In contrast, the FR-IQA approach hardly captures high-level semantics, and its labels ignore the impact of content perception on quality assessment.
V Experiment
This section first presents the experimental setting covering databases, protocols, and implementation details. After that, the experimental results are compared to state-of-the-art BIQA methods on ten IQA databases. Finally, we conducted comprehensive ablation studies to further analyze the effectiveness of our approach. thoroughly analyze the effectiveness of our approach.
V-A Experimental Settings
V-A1 Databases and protocols

### v2

@@ II-B2 Semantic-Aligned Quality Transfer (SAQT)-IQA dataset
Figure 4: The framework of the proposed GlintIQA. The input image is processed by VGFE and CLFE, followed by feature alignment and progressive integration using interactively stacked CWSA and SIEM. The resulting multi-grained representation is then fed to an MLP for quality prediction.
@@ III The Proposed GlintIQA
As shown in Fig.4, the proposed GlintIQA operates in three stages: global and local feature extraction (section III-A) via hybrid structures, i.e., ViT-based global feature extractor (VGFE) and CNNs-based local feature extractor (CLFE); global and local feature integration (section III-B) using channel-wise self-attention (CWSA) and spatial interaction enhancement module (SIEM) for multi-grained representations; and image quality prediction (section III-C) with MLP to predict the perceptual quality score.
@@ III-A Global and Local Feature Extraction
III-A1 VGFE
@@
$\textbf{f}_{g}^{e}=\psi_{\tiny f}\left(\mathcal{C}\left(\textbf{x};~{}3,d,k\right)\right)^{\text{T}},$ (4)
where the notation $\mathcal{C}\left(\textbf{x};3,d,k\right)$ denotes a 2D convolutional layer that takes an input x with 3 channels and transforms it into an output feature map with $d$ dimensions, using a kernel size of $k\times k$ and a stride of $k$. The function ${{\Large\psi}}_{\tiny f}$ represents the flattening operation that collapses the output of the convolutional layer along the spatial dimensions, yielding $\textbf{f}_{g}^{e}\in\mathbb{R}^{d\times\frac{HW}{k^{2}}}$. Subsequently, $\textbf{f}_{g}^{e}$ is transposed to obtain features within $\mathbb{R}^{\frac{HW}{k^{2}}\times d}$, where each patch of size $k\times k$ is represented by a vector of length $d$, referred to as a token. For simplicity, we have omitted the description of the class token and positional embedding used in the original ViT architecture.
@@
(a) The structure of CWSA
(b) The structure of SIEM
Figure 6: Structures of (a) channel-wise Self-attention (CWSA) and (b) Spatial Interaction Enhancement Module (SIEM).
$\textbf{f}_{g}^{i}=\mathcal{E}_{g}^{i}\left(\textbf{x};~{}\omega_{g}^{i}\right),$ (5)
Although the patch embedding process, as shown in Eq. 4, enables subsequent models to capture global context, it may result in the loss of fine-grained details within patches. As shown in Fig. 5(a), patch embedding discretizes the input image into a sequence of visual tokens, significantly reducing spatial dimensions in a single step at the model input. This discretization inevitably leads to a loss of fine-grained details and local spatial information, as the individual tokens correspond to low-resolution patches rather than individual pixels. To address this limitation, we exploit a local feature extractor consisting of a series of convolutional layers, as depicted in Fig. 5(b). This extractor uses small convolutional kernels and strides to progressively reduce the dimensions of the image, which can preserve certain local image details in the extracted features. Moreover, the integrated CNNs-based structure can mitigate the inherent lack of inductive bias in ViTs. By leveraging the strengths of both CNNs and ViTs, the hybrid model effectively captures both global and local features, ensuring a more accurate and robust representation of image quality.
III-A2 CLFE
Local features (e.g., edges and textures) reside within small image regions and provide valuable insights into the nature and severity of local distortions. To extract these features, the CLFE utilizes three residual blocks from ResNet50 [36], denoted as $\mathcal{E}_{l}^{j}$ with parameters $\omega_{l}^{j}$, where $j$ represents the index of the blocks. Each block yields features at distinct scales, downsampled by factors of $2^{j+1}$ relative to the original image.
@@ III-A2 CLFE
$\textbf{f}_{l}^{j}={\Large\psi}_{\tiny f}\left(\mathcal{C}\left({\mathcal{E}}_{l}^{j}\left(\textbf{x};\omega_{l}^{j}\right);~{}c_{j},d,k/2^{j+1}\right)\right)^{\text{T}},$ (6)
@@
$\textbf{f}_{l}=\mathcal{P}\left(\mathcal{M}_{\text{CWSA}}\left(\textbf{f}_{l}^{j-1}\oplus\textbf{f}_{l}^{j}\oplus\textbf{f}_{l}^{j+1}\right);~{}3d,d\right),$ (7)
where $\textbf{f}_{l}\in\mathbb{R}^{\frac{HW}{K^{2}}\times d}$ and $\oplus$ denotes a concatenation operation, and $\mathcal{P}(\cdot;3d,d)$ denotes a linear projection layer with input and output dimensions of $3d$ and $d$, respectively. The CWSA module, denoted as $\mathcal{M}_{\text{CWSA}}$, is described in detail in the following subsection.
III-B Global and Local Feature Integration
Integrating local and global features presents several challenges. Local features from CNNs and global features from ViTs operate at different scales, capturing fine-grained and coarse-grained details, respectively. A naive fusion of these multi-scale features risks diluting their distinct advantages and fails to capture the intricate dependencies between them. Additionally, local and global features contribute unequally across channels, with some local channels capturing critical texture details and others emphasizing high-level structural information. Thus, a more effective fusion mechanism is needed to integrate these multi-scale features while preserving their complementary nature. A dynamic, adaptive weighting mechanism is crucial to prioritize the most informative channels, ensuring a balanced and efficient fusion of local and global representations.
@@ III-B1 Progressive Feature Integration
Direct fusion of multi-granular features risks losing their complementary strengths by merging fine-grained and coarse-grained information simultaneously. To address this, a progressive integration strategy is employed, gradually merging features across layers. This approach allows the model to refine both local and global representations, optimizing them step-by-step across channels and spatial dimensions for more effective feature integration. Specifically, we start by concatenating the fine-grained feature $\textbf{f}_{l}$ with the coarse-grained global feature at the initial scale. Then, a stacked combination of CWSA and a SIEM is used to progressively fuse features in a hierarchical manner. The feature integration process can be concisely represented as:
@@
$\textbf{z}_{i}&=\left(\textbf{f}_{i}\oplus\textbf{f}_{g}^{i}\right)^{\text{T}};\\ \hat{\textbf{z}}_{i}&=\mathcal{M}_{\text{CWSA}}\left(\textbf{z}_{i}\right);\\ \textbf{f}_{i+1}&=\mathcal{M}_{\text{SIEM}}\left(\hat{\textbf{z}}_{i}\right).\end{aligned}\right.$ (8)
@@
For $i=1$, $\mathbf{f}_{i}$ represents the output feature of the CLFE, i.e., $\mathbf{f}_{l}$. For other values of $i$, $\mathbf{f}_{i}$ is the output feature of the SIEM, denoted as $\mathcal{M}_{\text{SIEM}}$.
III-B2 CWSA
Given the varying contributions of different channels in local and global features, CWSA is designed to dynamically learn and adjust the relative importance of each channel. This ensures that channels containing critical local texture details or key global structural information are appropriately emphasized. Using CWSA, the model computes the relevance of each channel to the task, reinforcing informative channels while down-weighting redundant ones. This adaptive weighting mechanism addresses the challenge of channel-wise importance disparity, enabling a more effective and balanced fusion process. As illustrated in Fig.6(a), CWSA concatenates two inputs to obtain the feature $\textbf{z}_{i}$ as shown in Eq.8. Subsequently, the feature is transposed and subjected to a self-attention computation through the following equations:
Q $=\mathcal{P}^{Q}(\textbf{z}_{i};d,d),\textbf{K}=\mathcal{P}^{K}(\textbf{z}_{i};d,d),\textbf{V}=\mathcal{P}^{V}(\textbf{z}_{i};d,d),$ (9)
$\hat{\textbf{z}}_{i}$ $=\text{Softmax}(\text{{Q}}\cdot\text{{K}}^{\text{T}}/\sqrt{d})\cdot\text{{V}}+\textbf{z}_{i},$ (10)
where $\textbf{z}_{i}$ is first projected into three distinct representations (query Q, key K, and value V) through separate linear projection layers $\mathcal{P}^{Q}$, $\mathcal{P}^{K}$, and $\mathcal{P}^{V}$. The channel-wise significance is then quantified through softmax scores applied to the scaled dot product of queries and keys, normalized by $\sqrt{d}$ ($d$ being the dimension). The final CWSA output $\hat{\textbf{z}}_{i}$ is obtained by combining the attention-weighted values with the residual connection of the input feature $\textbf{z}_{i}$, where $\text{Softmax}(\cdot)$ normalizes the values into a probability distribution over the output dimensions.
III-B3 SIEM
SIEM follows channel-wise attention to enhance spatial relationships between local and global features. By reshaping and applying convolutional operations, SIEM fosters deeper interactions across multiple scales, effectively capturing spatial dependencies and complementing CWSA to produce a representation that is both channel-optimized and spatially robust. The module operates on the input tensor $\hat{\textbf{z}}_{i}$, as illustrated in Fig.6(b). The computational process is expressed as:
$\textbf{f}_{i+1}=\psi_{f}\left(\mathcal{C}\left(\varphi_{r}\left(\hat{\textbf{z}}_{i}\right);~{}2d,d,3\right)\right),$ (11)
@@ III-B3 SIEM
where the notation ${\Large\varphi}_{\tiny r}$ indicates the operation of reshaping the token vectors into 2D feature maps. The convolutional layer $\mathcal{C}(\cdot;2d,d,3)$ employs a $3\times 3$ filter with a stride of 1. Finally, the output of the convolutional layer is reshaped back to a flattened tensor, aligning it with the original input dimensions.
TABLE I: The Summary of IQA Databases for NR-IQA Models.
Database Database Ref. Image Dist. Type Dist. Image Image
Type Name Count Count Count Resolution
Synthetic LIVE [28] 29 5 779 480$\times$720 $\sim$ 768$\times$512
CSIQ [33] 30 6 866 512$\times$512
TID2013 [29] 25 24 3,000 384$\times$512
KADID-10k [30] 81 25 10,125 384$\times$512
LIVE-MD [34] 15 2 405 720$\times$1280
Authentic BID [37] - - 585 960$\times$1280 $\sim$ 1704$\times$2272
LIVEC [38] - - 1,162 500$\times$500 $\sim$ 640$\times$960
KonIQ-10k [39] - - 10,073 384$\times$512/768$\times$1024
SPAQ [40] - - 11,125 1080$\times$1440$\sim$6556$\times$3744
LIVEFB [41] - - 39,810 160$\times$186 $\sim$ 660$\times$1200
III-C Image Quality Prediction
@@
The quality score prediction is performed via an MLP architecture with two linear projection layers. The feature map f from the SIEM module is spatially averaged to obtain a $d$-dimensional vector $\hat{\textbf{f}}\in\mathbb{R}^{1\times d}$, forming a compact quality representation. This vector is then processed through two linear projections with GELU activation and dropout regularization to generate the final quality score $s$:
$s=\mathcal{P}\left(\mathcal{P}(\hat{\textbf{f}};~{}d,d);d,1\right).$ (12)
@@
where $\mathcal{P}$ denotes a linear projection layer. For simplicity, the specific details of the activation function and dropout layer are omitted from the above equation. Ultimately, $l_{1}$ loss function is utilized to enable end-to-end model parameter updates.
IV Experiment
This section begins by outlining the experimental setup (section IV-A). Subsequently, we compare the experimental results against state-of-the-art NR-IQA methods across ten IQA databases, encompassing both single-dataset (section IV-B) and cross-dataset (section IV-C) evaluations. Finally, we perform comprehensive ablation studies to further assess the effectiveness of our approach (section IV-D).
Figure 7: The pre-training results of the proposed SAQT-IQA dataset.
IV-A Experimental Settings
IV-A1 Databases and protocols